# Federated Large Language Models: Current Progress and Future Directions

### v1

@@ 4.4 Federated Inference on Local Data
Another promising direction is the development of techniques for federated inference, where models trained in a federated manner are used to make predictions directly on local data. This approach would help mitigate privacy concerns while allowing for real-time, on-device inference. However, optimizing the inference process to minimize latency and computational overhead remains challenging. Prior works such as [40] adapt pre-trained models for federated contexts during inference using parameter-efficient learning to minimize the computational load. [44] has also explored aligning local representations with global models to enhance the accuracy and responsiveness in federated inference. While these works focus on smaller models, future work may develop new federated inference techniques for LLMs, addressing the trade-offs between model complexity, real-time performance, and resource efficiency, which are significant for large-scale models.
@@ 4.5 LLMs for Federated Learning
As we discussed above in Section 1, concerns about data privacy and data scarcity can severely limit the effectiveness of models in FL scenarios  [50, 19, 7], especially when sensitive or limited data are involved, such as in the healthcare or financial sectors [38, 92]. Large Language Models propose a promising solution by generating synthetic data, which can significantly enrich the original training sets [65, 34, 69]. This synthetic data, being reflective of diverse real-world scenarios, helps alleviate the issue of data scarcity and introduces additional diversity, thus enhancing the model’s generalization capabilities and reducing the likelihood of overfitting [20, 15, 76].
@@
Furthermore, LLMs offer a solid foundation for improving FL performance [88, 91] through advanced capabilities such as knowledge distillation [32, 83] and prompt engineering [98]. These models provide a starting point for local fine-tuning on clients’ data, which enables faster convergence and potentially higher model accuracy without the extensive need for ground-up model training [45, 9]. In addition, by leveraging the rich data understanding inherent in LLMs, suboptimal performance due to limited or imbalanced data can be substantially mitigated, thus improving efficiency and effectiveness across various FL deployments [91].
@@
However, the integration of LLMs into FL also brings forth complex challenges concerning law compliance and responsible technology usage [59, 16, 103, 66]. The generation of synthetic data by LLMs, while beneficial, must be carefully managed to avoid potential privacy infringements and intellectual property violations. Ensuring that synthetic data does not too closely resemble the original training sets is crucial to mitigate the risks of revealing sensitive information. These legal and ethical considerations necessitate a framework that not only advances the technological capabilities of FL but also aligns with stringent regulatory standards to ensure the responsible deployment of these powerful models in sensitive domains.

### v2

@@ 4.4 Federated AI Agents
Federated Inference on Local Data One promising direction is the development of techniques for federated inference, where models trained in a federated manner are used to make predictions directly on local data. This approach would help mitigate privacy concerns while allowing for real-time, on-device inference. However, optimizing the inference process to minimize latency and computational overhead remains challenging. Prior works such as [liu2022federated] adapt pre-trained models for federated contexts during inference using parameter-efficient learning to minimize the computational load. [ma2022federated] has also explored aligning local representations with global models to enhance the accuracy and responsiveness in federated inference. While these works focus on smaller models, future work may develop new federated inference techniques for LLMs, addressing the trade-offs between model complexity, real-time performance, and resource efficiency, which are significant for large-scale models.
Federated Agents across Clients Beyond on-device inference, an emerging direction is to treat each client as an agent, an LLM-driven entity that observes local environments, calls tools, and participates in federated optimization. In this view, FedLLM systems evolve into distributed populations of heterogeneous agents with local capabilities (e.g., private RAG or workflow execution) that align toward shared global objectives under privacy and resource constraints [jing2025federated, krouka2021communication].
Recent work explicitly explores LLM agents under FL. FICAL [wu2024federated] enables communication-efficient coordination by exchanging compact knowledge compendiums instead of parameters, while also teaching agents tool-use via RAG modules. FedAgent [sun2025fedagent] demonstrates that decentralized RL training can optimize agent policies across clients without sharing trajectories, achieving near-centralized performance while preserving data locality.
FMARL research provides key algorithmic foundations, such as federated multi-agent deep RL [jing2025federated] and heterogeneity-aware approaches like FedMRL [sahoo2024fedmrl], which support client-specific policies and robust aggregation across non-IID environments. These ideas transfer naturally to FedLLM agents that must collaborate across organizations with divergent goals.
At the systems level, frameworks such as Academy [pauloski2025empowering] deploy stateful LLM-based agents across federated research infrastructure, enabling distributed execution while preserving institutional autonomy. In federated RAG settings, each client acts as a local retrieval agent that queries its private corpus and contributes only minimal retrieval signals (top-k results or relevance scores) to the global generation process, avoiding raw data exchange [wang2024feb4rag].
Open challenges remain in coordination and credit assignment, communication efficiency, personalization with safety, and standardized evaluation. Addressing these issues will be crucial for transforming FedLLM from distributed model training into large-scale, privacy-preserving multi-agent systems.
@@ 4.5 LLMs for Federated Learning
As we discussed above in Section 1, concerns about data privacy and data scarcity can severely limit the effectiveness of models in FL scenarios  [FedAffect, gao2022survey, chen2022practical], especially when sensitive or limited data are involved, such as in the healthcare or financial sectors [liu2023efficient, zhang2022dense]. Large Language Models propose a promising solution by generating synthetic data, which can significantly enrich the original training sets [wang2022self, li2023synthetic, wei2023simple]. This synthetic data, being reflective of diverse real-world scenarios, helps alleviate the issue of data scarcity and introduces additional diversity, thus enhancing the model’s generalization capabilities and reducing the likelihood of overfitting [goetz2020federated, dankar2022using, wu2024prompt].
@@
Furthermore, LLMs offer a solid foundation for improving FL performance [yu2023federated, zhang2024mllm] through advanced capabilities such as knowledge distillation [li2019fedmd, yang2023knowledge] and prompt engineering [zhao2023fedprompt]. These models provide a starting point for local fine-tuning on clients’ data, which enables faster convergence and potentially higher model accuracy without the extensive need for ground-up model training [nguyen2022begin, chen2022importance]. In addition, by leveraging the rich data understanding inherent in LLMs, suboptimal performance due to limited or imbalanced data can be substantially mitigated, thus improving efficiency and effectiveness across various FL deployments [zhang2024mllm].
@@
However, the integration of LLMs into FL also brings forth complex challenges concerning law compliance and responsible technology usage [wachter2024large, darji2024challenges, zhong2020does, wang2023where]. The generation of synthetic data by LLMs, while beneficial, must be carefully managed to avoid potential privacy infringements and intellectual property violations. Ensuring that synthetic data does not too closely resemble the original training sets is crucial to mitigate the risks of revealing sensitive information. These legal and ethical considerations necessitate a framework that not only advances the technological capabilities of FL but also aligns with stringent regulatory standards to ensure the responsible deployment of these powerful models in sensitive domains.